OpenStudy (konradzuse):

Find the area of the region. Inside r = 2a cos(θ) and outside r = a

OpenStudy (anonymous):

first find the points of intersections solve a=2acos(theta) let me know what you got.

OpenStudy (konradzuse):

hmmm

OpenStudy (konradzuse):

how to find poi?

OpenStudy (anonymous):

solve this equation \[\Large a=2acos(\theta)\]

OpenStudy (konradzuse):

solve for theta? a/a = 1 = 2cos(theta) arccos(1/2) = theta?

OpenStudy (anonymous):

yes correct

OpenStudy (konradzuse):

now what do I do?

OpenStudy (konradzuse):

or is that the answer........?

OpenStudy (anonymous):

take your calculator and find arccos(1/2)

OpenStudy (konradzuse):

pi/3

OpenStudy (anonymous):

another value as well. because the graph is intersecting in 4th Quadrant as well.

OpenStudy (konradzuse):

5pi/3?

OpenStudy (anonymous):

do you think 2pi/3 is in 4th quadrant :P

OpenStudy (konradzuse):

I'm really bad with this stuff sorry.. Some of my basic math was destroyed by horrid teachers...

OpenStudy (anonymous):

ok the other angle is \[\Large 2\pi-\pi/3=5\pi/3\] sorry i did not look above you had found it it.

OpenStudy (konradzuse):

:)

OpenStudy (anonymous):

ok now the range is in our hand s up integral \[\Large Area=\int\limits_{a}^{b}\frac{1}{2}r^2d \theta\]

OpenStudy (konradzuse):

int from pi/3 to 5pi/3 of 1/2 r^2d theta huh?

OpenStudy (konradzuse):

which would equal r^3/6 from those points :P.

OpenStudy (anonymous):

\[\Large \int\limits_{\frac{5\pi}{3}}^{\frac{\pi}{3}}\frac{1}{2}(2acos(\theta))^2-a^2)d \theta\] can you do this now ?

OpenStudy (konradzuse):

5pi/3 is on the bottom.........?

OpenStudy (konradzuse):

@sami-21

OpenStudy (konradzuse):

also how do we use the a's in this? Confused stilll :(.

OpenStudy (anonymous):

yes it is in the bottom. also treat a as a constant.

OpenStudy (konradzuse):

I mean we solved for it, but you want me to use it in that form?

OpenStudy (anonymous):

use it in that form . just ntegrate the above integral and you have the required area.

OpenStudy (konradzuse):

well can we cancel them? It looks like we can get rid of the 1/2 with the 2 in front.

OpenStudy (anonymous):

2 is under the square (2acos(theta))^2

OpenStudy (konradzuse):

true.......... hmm....

OpenStudy (konradzuse):

you're missing an open paren up there btw....

OpenStudy (konradzuse):

Maybe that'sd why I'm so konfused....

OpenStudy (anonymous):

yes i am missing one paren. :P

OpenStudy (konradzuse):

Idk if I've ever seen an integral with 2 variables...

OpenStudy (konradzuse):

theta would be x.... but you said a is a "constant?" I mean but it could be anything... right?

OpenStudy (konradzuse):

:O

OpenStudy (konradzuse):

@sami-21 done?

OpenStudy (anonymous):

there is only obe variable thetha. let me explain little bit.

OpenStudy (konradzuse):

ok ty :)

OpenStudy (anonymous):

\[\Large \frac{1}{2} \int\limits_{\frac{5\pi}{3}}^{\frac{\pi}{3}}[4a^2\cos^2(\theta)-a^2)d \theta\] or \[\Large \frac{a^2}{2} \int\limits\limits_{\frac{5\pi}{3}}^{\frac{\pi}{3}}[(4\cos^2(\theta)-1)d \theta]\] i hope now you can do it.

OpenStudy (konradzuse):

hmm

OpenStudy (konradzuse):

:)

OpenStudy (konradzuse):

a^2/2[(pi/3+sin(2pi/3)) -(5pi/3 + sin(10pi/3))] ????

OpenStudy (konradzuse):

which = sqrt(3)-(4 π)/3

OpenStudy (konradzuse):

with a^2/2 also there.

OpenStudy (anonymous):

you have missed second term while integrating there should be second term as well when you expand integral sign. \[\Large \int\limits_{\frac{5\pi}{3}}^{\frac{\pi}{3}} 1d \theta\]

OpenStudy (konradzuse):

:(?